OpenStudy (fuzzyball):

FAN AND MEDAL A college has a student to teacher ratio of 30 to 2. If there are 186 teachers at the college, how many students attend the college? A. 2,790 B. 3,906 C. 1,116 D. 2,232

OpenStudy (fortytherapper):

Setting up a proportion should work. Let's first read and see which makes sense. It says the ratio is 30 to 2, so would that mean 30 students per 2 teachers or 30 teachers per 2 students?

OpenStudy (fuzzyball):

30 students per 2 teachers

OpenStudy (fortytherapper):

Right. So we have 2 teachers to 30 students And we have 186 teachers to ? students We can set it up now \[\frac{ 2 }{ 30 } = \frac{ 186 }{ x }\] Do you know how to cross multiply and divide?

OpenStudy (fuzzyball):

no, maybe would i cross multiply or divide in the problem?

OpenStudy (fortytherapper):

Yes So basically, you would get the top number and make a diagonal line to the other value So for 2. You create that line diagonally and it goes to x. For 30, you do the same thing and it goes to 186 (It also formed a x-cross) So now you multiply the numbers you crossed, so \[2x = (186)(30)\] Now solve for (186)(30), what do you get?

OpenStudy (fuzzyball):

5,580

OpenStudy (fortytherapper):

Yes! So now it's \[2x = 5580\] But we just want our answer as an x, not 2x. Since it's a 2 multiplied by an X, we have to do the opposite of that to get it to x, and we have to do it to both sides

OpenStudy (fuzzyball):

2,790

OpenStudy (fortytherapper):

Correct, good work

OpenStudy (fuzzyball):

tysm!!

OpenStudy (fortytherapper):

No problem!

OpenStudy (fuzzyball):

can you help me with 1 more ill open in new thing so i can medal u

OpenStudy (fortytherapper):

Sure